# New results for the Cahn-Hilliard equation with non-degenerate mobility: well-posedness and longtime behavior

### v1

@@ 1. Introduction and main results
Here, $\Omega$ is a bounded domain in $\mathbb{R}^{d}$, $d\leq 3$, and n is the outward normal vector on $\partial\Omega$. The state variable $\varphi$ represents the difference of phase concentrations, and takes values in the interval $[-1,1]$, with $\varphi=\pm 1$ being the pure phases. The function $b:[-1,1]\to\mathbb{R}$ is the so-called Onsager mobility, which measures the strength of the diffusion. The mobility can assume one the following forms
@@
    Let $\varphi_{0}\in H^{1}(\Omega)\cap L^{\infty}(\Omega)$ be such that $\left\|\varphi_{0}\right\|_{L^{\infty}}\leq 1$ and $\overline{\varphi_{0}}\in(-1,1)$. Then, there exists a global weak solution $\left(\varphi,\mu\right)$ to problem (1.1a)-(1.1b) and (1.2) satisfying
    $\begin{split}&\varphi\in L^{\infty}(0,T;H^{1}(\Omega))\cap L^{2}(0,T;H^{2}(\Omega))\cap H^{1}(0,T;H_{(0)}^{-1}(\Omega)),\\ &\mu\in L^{2}(0,T;H^{1}(\Omega)),\quad\Psi^{\prime}(\varphi)\in L^{2}(0,T;L^{2}(\Omega)),\end{split}$ (1.8)
@@
    $\begin{split}&\varphi\in L^{\infty}(0,T_{\ast};H^{2}(\Omega))\cap H^{1}(0,T_{\ast};H^{1}(\Omega))\cap W^{1,\infty}(0,T_{\ast};H_{(0)}^{-1}(\Omega)),\\ &\mu\in L^{\infty}(0,T_{\ast};H^{1}(\Omega))\cap L^{2}(0,T_{\ast};H^{2}(\Omega)),\end{split}$ (1.12)
@@
The results presented in Theorem 1.1 exhibit some faults. Firstly, the condition in (1.11) on the time derivative $\partial_{t}\varphi$ is never reached by weak solutions (cf. (1.8)). Thus, no uniqueness result regarding the weak solutions is known, leaving this issue an open question in the Cahn-Hilliard theory. Secondly, concerning the global existence of strong solutions in the two dimensional case (cf. Theorem 1.1 - (3)), we point out that the proof of [5, Corollary 2.1] contains a slight but crucial error. Precisely, in the so-called second energy estimates involving
@@
The goal of this contribution is to address the gaps left open by Theorem 1.1 concerning the well-posedness theory of weak solutions to the Cahn-Hilliard equation with non-degenerate mobility in two dimensions. First, we will prove the uniqueness of weak solutions. Then, we will show that the weak solutions exhibit a global smoothness effect at positive times (propagation of regularity). As a consequence, we will also prove that the weak solution converges towards a single steady state as $t\to\infty$. More precisely, our result is the following:
@@
    $\partial_{t}\varphi\in L^{2}(0,\infty;H_{(0)}^{-1}(\Omega)),\quad F^{\prime}(\varphi)\in L_{\operatorname{uloc}}(0,T;L^{p}(\Omega)),$ (1.16)
@@
    Uniqueness of weak solutions. Suppose that $b\in C^{2}([-1,1])$. Let $\varphi^{1},\varphi^{2}$ be two weak solutions originating from the initial conditions $\varphi^{1}_{0},\varphi^{2}_{0}$, respectively, such that $\overline{\varphi^{1}_{0}}=\overline{\varphi^{2}_{0}}$. Then, for any $T>0$, there exists a positive constant $C$ such that
    $\left\|\varphi_{1}(t)-\varphi_{2}(t)\right\|_{H_{(0)}^{-1}(\Omega)}\leq C\left\|\varphi_{0}^{1}-\varphi_{0}^{2}\right\|_{H_{(0)}^{-1}(\Omega)}\!,\quad\forall\,t\in[0,T].$ (1.20)
    The constant $C$ only depends on the parameters of the system, the final time $T$, and the initial free energies $E(\varphi_{0}^{1})$ and $E(\varphi_{0}^{2})$. In particular, the weak solution is unique.
@@
    Convergence to a stationary point. If $b\in C^{2}([-1,1])$, the solution $\varphi(t)$ converges to $\varphi_{\infty}$ in $W^{2-\varepsilon,p}(\Omega)$ as $t\rightarrow\infty$, for any $\varepsilon>0$ and any $2\leq p<\infty$, where $\varphi_{\infty}\in W^{2,p}(\Omega)$ is a solution to the stationary Cahn-Hilliard equation
@@
  • We observe that the validity of the energy equality (1.19) for weak solutions was already shown in [41, Lemma 2.4].
@@
$\|\sqrt{b(\varphi_{1})}\nabla\mathcal{G}_{\varphi_{1}}(\varphi_{1}-\varphi_{2})\|_{L^{2}(\Omega)},$ (1.26)
where the operator $\mathcal{G}_{\varphi_{1}}$ is defined in (2.9). In order to derive a suitable differential inequality for this quantity, we will address two crucial steps. Firstly, we need to rigorously justify the integration by parts formula (3.16) involving the time derivative of (1.26). Secondly, we will provide a novel control of the non-linear terms in (3.18)-(3.20). Both steps rely on the combination of two analytical tools: an elliptic regularity theory for a Neumann problem in divergence form with concentration-dependent coefficient (cf. (2.8)), and the regularity $L_{\operatorname{uloc}}^{4}([0,\infty);H^{2}(\Omega))$ for weak solutions, firstly discovered in [27]. Regarding the propagation of regularity, we will first revise the approach in [5]. By carefully examining the dependence of the constants appearing in the proof, our analysis shows that the correct differential inequality for $\Lambda$ reads as follows
@@
for some positive constant $C$ independent of $\rho$, which differs from (1.13). Due to the factor $\frac{1}{\rho}$, the argument in [5] do not lead to a global bound for arbitrarily large initial datum. To overcome this issue, we will exploit a recent Grownall-type argument devised in [36, Lemma 2.1] (see Lemma 4.2 below). However, a direct application of this result leads to a global bound which grows double exponentially in time (4.25). Then, we propose a new approach to estimate the nonlinear term in (4.8). Once again, it relies on a careful use of the elliptic regularity theory for (2.8). Thanks to the regularity $\varphi\in L_{\operatorname{uloc}}^{4}([0,\infty);H^{2}(\Omega))$, the final differential inequality (4.27) fits into the classical framework of the uniform Gronwall lemma, thus implying the desired global estimate independent of time. Finally, the proof of the convergence of the weak solution towards a steady state of the Cahn-Hilliard equation is performed by following the well established method in the constant mobility case (e.g., [4]). This hinges upon the regularity of the weak solutions, the presence of a strict Lyapunov function, as well as the Łojasiewicz-Simon inequality.
@@ 2.1. Notation and function spaces
$\left\|f\right\|_{L^{2}(\Omega)}\leq C_{P}\left(\left\|\nabla f\right\|_{L^{2}(\Omega)}+|\overline{f}|\right)\!,\quad\forall\,f\in H^{1}(\Omega).$ (2.3)
@@
The solution operator [^1] Despite the recent literature makes use of the symbol $\mathcal{N}$, we keep the notation used in [5]. $\mathcal{G}:H^{-1}_{(0)}(\Omega)\to H^{1}_{(0)}(\Omega)$ is defined as follows: for every $f\in H^{-1}_{(0)}(\Omega)$, $\mathcal{G}f\in H^{1}_{(0)}(\Omega)$ is the unique function satisfying
@@
where $F$ is the convex part of $\Psi$ as in (1.6). The functional $E_{0}$ is proper, convex and lower-semicontinuous. Owing to [40, Lemma 4.1], we deduce that $t\mapsto E_{0}(\varphi(t))$ is absolutely continuous on $[0,\infty)$ and
@@
for almost everywhere on $(0,\infty)$. As a consequence, for any $T>0$, due to the boundedness of $F$ and $\varphi\in C([0,T];L^{2}(\Omega))$, the Lebesgue theorem entails that $\int_{\Omega}F(\varphi(\cdot))\ {\rm d}x\in C([0,T]),$ which in turn gives $\varphi\in C([0,T];H^{1}(\Omega))$. Now, taking $v=\mu$ in (1.18a) and exploiting the standard chain rule in $L^{2}(0,T;H^{1}(\Omega))\cap H^{1}(0,T;H^{1}(\Omega)^{\prime})$, we obtain
@@
for some constant $C$ depending on $E(\varphi_{0})$ and $m$. The latter implies $\nabla\mu\in L^{2}(0,\infty;L^{2}(\Omega))$. Since by (3.1) and (2.10), we have
@@
Next, multiplying (1.18b) by $-\Delta\varphi$ and exploiting the boundary condition on $\varphi$, we get
@@
It is well known that the second term on the left hand side is non-negative by the convexity of $F$ (see, e.g., [13, Appendix A]). Then, by (3.5), it follows that
@@
where the positive constant $C$ depends on $E(\varphi_{0})$ and $m$. Since $|\overline{\mu}|=\overline{\Psi^{\prime}(\varphi)},$ (3.12) and the above inequality imply
@@
$\begin{cases}-\Delta\varphi+F^{\prime}(\varphi)=\mu^{*},\quad&\text{in }\Omega,\\ \partial_{\textbf{{n}}}\varphi=0,\quad&\text{on }\partial\Omega,\end{cases}$
@@
$\mu_{i}=-\mathcal{G}_{\varphi_{i}}\partial_{t}\varphi_{i}+\overline{\Psi^{\prime}(\varphi_{i})},\quad\text{ in }\Omega,\ \text{a.e. in }(0,\infty),\ i=1,2.$ (3.15)
We define $\Phi=\varphi_{1}-\varphi_{2}$, which solves
@@
where $\mu_{1}$ and $\mu_{2}$ are the chemical potentials corresponding to $\varphi_{1}$ and $\varphi_{2}$, respectively. We preliminary notice that $\overline{\Phi}\equiv 0$ by the conservation of mass. Multiplying the above equation by $\Phi$ and integrating over $\Omega$, we find
@@
Thanks to (3.15), we rewrite the above as follows
@@
almost everywhere in $(0,\infty)$. Since $\varphi_{1}\in L^{\infty}(0,T;H^{1}(\Omega))\cap L^{4}(0,T;H^{2}(\Omega))\cap H^{1}(0,T;H^{-1}_{(0)}(\Omega))$, for any $T>0$, it follows by the elliptic regularity theory that
$\varphi_{1}^{\varepsilon}\in L^{\infty}(0,T;H^{3}(\Omega))\cap H^{1}(0,T;H^{1}_{(0)}(\Omega)),\quad\forall\,\varepsilon\in(0,1),\ \forall\,T>0,$ (3.22)
and $\partial_{\textbf{{n}}}\varphi_{1}^{\varepsilon}=0$ on $\partial\Omega$ almost everywhere in $(0,\infty)$, for any $\varepsilon\in(0,1)$. In addition, recalling that $\varphi_{1}\in L^{\infty}(\Omega\times(0,T))$ with $|\varphi_{1}|<1$ almost everywhere in $\Omega\times(0,T)$, we also have
@@
$\int_{0}^{T}\|\varphi_{1}^{\varepsilon}(s)\|_{L^{2}(\Omega)}^{4}+\|\nabla\varphi_{1}^{\varepsilon}(s)\|_{L^{2}(\Omega)}^{4}\,{\rm d}s$ $\leq\int_{0}^{T}\|\varphi_{1}(s)\|_{L^{2}(\Omega)}^{4}+\|\nabla\varphi_{1}(s)\|_{L^{2}(\Omega)}^{4}\,{\rm d}s,$ (3.24)
@@
Owing to the regularity of $\varphi_{1}^{\varepsilon}$, we compute the gradient of (3.21) and test by $-\nabla\Delta\varphi_{1}^{\varepsilon}$. Exploiting the Neumann boundary condition on $\varphi_{1}^{\varepsilon}$ and $\varphi_{1}$, we find
@@
$\varphi_{1}^{\varepsilon}\to\varphi_{1}\ \text{strongly in }L^{4}(0,T;H^{2}(\Omega)),\quad\partial_{t}\varphi_{1}^{\varepsilon}\to\partial_{t}\varphi_{1}\ \text{strongly in }L^{2}(0,T;H^{-1}_{(0)}(\Omega)).$ (3.28)
@@
$\|\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\|_{L^{\infty}(0,T;H^{1}_{(0)}(\Omega))}\leq C,$ (3.30)
@@
$\|\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\|_{H^{2}(\Omega)}\leq C\left(1+\|\varphi_{1}^{\varepsilon}\|_{H^{2}(\Omega)}\right)\!,$
@@
$\int_{0}^{T}\|\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi(s)\|_{H^{2}(\Omega)}^{4}\,{\rm d}s\leq C,$ (3.31)
@@
$\|\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\|_{W^{2,4}(\Omega)}\leq C\left(1+\|\varphi_{1}^{\varepsilon}\|_{H^{2}(\Omega)}^{\frac{3}{4}}\|\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\|_{H^{2}(\Omega)}^{\frac{3}{4}}\right)\!.$ (3.32)
@@
$\int_{0}^{T}\|\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi(s)\|_{W^{2,4}(\Omega)}^{\frac{8}{3}}\,{\rm d}s\leq CT+C\int_{0}^{T}\|\varphi_{1}^{\varepsilon}(s)\|_{H^{2}(\Omega)}^{4}+\|\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi(s)\|_{H^{2}(\Omega)}^{4}\,{\rm d}s\leq C,$ (3.33)
@@
$\int_{0}^{T}\|\nabla\left(\mathcal{G}_{\varphi_{1}}f-\mathcal{G}_{\varphi_{1}^{\varepsilon}}f\right)\|_{L^{2}(\Omega)}^{2}\,{\rm d}s$ $\leq\frac{1}{b_{m}}\int_{0}^{T}\|\left(b(\varphi_{1}^{\varepsilon})-b(\varphi_{1})\right)\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}f\|_{L^{2}(\Omega)}^{2}\,{\rm d}s$
$\leq\frac{1}{b_{m}}\int_{0}^{T}\|b(\varphi_{1}^{\varepsilon})-b(\varphi_{1})\|_{L^{\infty}(\Omega)}^{2}\|\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}f\|_{L^{2}(\Omega)}^{2}\,{\rm d}s.$
We aim to use the Lebesgue convergence theorem on the right hand side. We observe from (2.10) that
$\|b(\varphi_{1}^{\varepsilon})-b(\varphi_{1})\|_{L^{\infty}(\Omega)}^{2}\|\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}f\|_{L^{2}(\Omega)}^{2}\leq\frac{4b_{M}}{b_{m}}\|\nabla\mathcal{G}f\|_{L^{2}(\Omega)}^{2},$
@@
$\|b(\varphi_{1}^{\varepsilon})-b(\varphi_{1})\|_{L^{\infty}(\Omega)}^{2}\leq C\|\varphi_{1}^{\varepsilon}-\varphi_{1}\|_{L^{\infty}(\Omega)}^{2}\leq C\|\varphi_{1}^{\varepsilon}-\varphi_{1}\|_{H^{2}(\Omega)}^{2}$
@@
$\|b(\varphi_{1}^{\varepsilon})-b(\varphi_{1})\|_{L^{\infty}(\Omega)}^{2}\|\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}f\|_{L^{2}(\Omega)}^{2}\leq C\|\varphi_{1}^{\varepsilon}-\varphi_{1}\|_{H^{2}(\Omega)}^{2}\|\nabla\mathcal{G}f\|_{L^{2}(\Omega)}^{2}\to 0$
@@
$\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}f\to\nabla\mathcal{G}_{\varphi_{1}}f\ \text{strongly in }L^{2}(0,T;L^{2}(\Omega)).$

### v2

@@ Abstract.
Keywords: Cahn-Hilliard equation, non-degenerate mobility, uniqueness, regularity, convergence to steady states
MSC 2010: 35K55, 35A02, 35B65, 35B40, 35Q82
@@ 1. Introduction and main results
Here, $\Omega$ is a bounded domain in $\mathbb{R}^{d}$, $d\leq 3$, and n is the outward normal vector on $\partial\Omega$. The state variable $\varphi$ represents the difference of phase concentrations in a binary mixture, and takes values in the interval $[-1,1]$, with $\varphi=\pm 1$ being the pure phases. The function $b:[-1,1]\to\mathbb{R}$ is the so-called Onsager mobility, which measures the strength of the diffusion. The mobility can assume one the following forms
@@
    Let $\varphi_{0}\in H^{1}(\Omega)\cap L^{\infty}(\Omega)$ be such that $\left\|\varphi_{0}\right\|_{L^{\infty}(\Omega)}\leq 1$ and $\overline{\varphi_{0}}=\frac{1}{|\Omega|}\int_{\Omega}\varphi_{0}(x)\,{\rm d}x\in(-1,1)$. Then, there exists a global weak solution $\left(\varphi,\mu\right)$ to problem (1.1a)-(1.1b) and (1.2) satisfying
    $\varphi\in L^{\infty}(0,T;H^{1}(\Omega))\cap L^{2}(0,T;H^{2}(\Omega)),\quad\partial_{t}\varphi\in L^{2}(0,T;H_{(0)}^{-1}(\Omega)),$ (1.8)
    $\mu\in L^{2}(0,T;H^{1}(\Omega)),\quad\Psi^{\prime}(\varphi)\in L^{2}(0,T;L^{2}(\Omega)),$ (1.9)
@@
    $\varphi\in L^{\infty}(0,T_{\ast};H^{2}(\Omega))\cap H^{1}(0,T_{\ast};H^{1}(\Omega)),\quad\partial_{t}\varphi\in L^{\infty}(0,T_{\ast};H_{(0)}^{-1}(\Omega)),$ (1.13)
    $\mu\in L^{\infty}(0,T_{\ast};H^{1}(\Omega))\cap L^{2}(0,T_{\ast};H^{2}(\Omega)),$ (1.14)
@@
The results presented in Theorem 1.1 exhibit some shortcomings. Firstly, the condition in (1.12) on the time derivative $\partial_{t}\varphi$ is never reached by weak solutions (cf. (1.8)-(1.9)). Thus, no uniqueness result regarding the weak solutions is known, leaving this issue an open question in the Cahn-Hilliard theory. Secondly, concerning the global existence of strong solutions in the two dimensional case (cf. Theorem 1.1 - (3)), we point out that the proof of [5, Corollary 2.1] contains a slight but crucial error. Precisely, in the so-called second energy estimates involving
@@
The goal of this contribution is to address the gaps left open by Theorem 1.1 concerning the well-posedness theory of weak solutions to the Cahn-Hilliard equation with non-degenerate mobility in two dimensions. First, we will prove the uniqueness of weak solutions. Then, we will show that the weak solutions exhibit a global smoothness effect at positive times (propagation of regularity). As a consequence, we will prove that any weak solution converges towards a single steady state as $t\to\infty$.
In order to rigorously state our achievements, we direct the reader to Section 2 for the necessary notation, in particular concerning the Bochner spaces $BC([0,\infty);X)$ and $L_{\rm uloc}^{p}([0,\infty);X)$ of functions that are respectively bounded and continuous in time, and locally uniformly integrable in time, with values in a Banach space $X$.
Our main result is as follows:
@@
    $\partial_{t}\varphi\in L^{2}(0,\infty;H_{(0)}^{-1}(\Omega)),\quad F^{\prime}(\varphi)\in L_{\rm uloc}^{2}([0,\infty);L^{p}(\Omega)),$ (1.18)
@@
    Uniqueness of weak solutions. Suppose that $b\in C^{2}([-1,1])$. Let $\varphi_{1},\varphi_{2}$ be two weak solutions originating from the initial conditions $\varphi_{1}^{0},\varphi_{2}^{0}$, respectively, such that $\overline{\varphi_{1}^{0}}=\overline{\varphi_{2}^{0}}$. Then, for any $T>0$, there exists a positive constant $C$ such that
    $\left\|\varphi_{1}(t)-\varphi_{2}(t)\right\|_{H_{(0)}^{-1}(\Omega)}\leq C\left\|\varphi_{1}^{0}-\varphi_{2}^{0}\right\|_{H_{(0)}^{-1}(\Omega)}\!,\quad\forall\,t\in[0,T].$ (1.22)
    The constant $C$ depends only on the parameters of the system, the final time $T$, and the initial free energies $E(\varphi_{1}^{0})$ and $E(\varphi_{2}^{0})$. In particular, the weak solution is unique.
@@
    Convergence to a stationary state. If $b\in C^{2}([-1,1])$, the solution $\varphi(t)$ converges to $\varphi_{\infty}$ in $W^{2-\varepsilon,p}(\Omega)$ as $t\rightarrow\infty$, for any $\varepsilon>0$ and any $2\leq p<\infty$, where $\varphi_{\infty}\in W^{2,p}(\Omega)$ is a solution to the stationary Cahn-Hilliard equation
@@
  • We observe that the validity of the energy equality (1.21) for weak solutions was already shown in [40, Lemma 2.4].
@@
$\left\|\sqrt{b(\varphi_{1})}\nabla\mathcal{G}_{\varphi_{1}}(\varphi_{1}-\varphi_{2})\right\|_{L^{2}(\Omega)},$ (1.28)
where the operator $\mathcal{G}_{\varphi_{1}}$ is defined in (2.9). In order to derive a suitable differential inequality for this quantity, we will address two crucial steps. Firstly, we need to rigorously justify the integration by parts formula (3.16) involving the time derivative of (1.28). Secondly, we will provide a novel control of the non-linear terms in (3.18)-(3.20). Both steps rely on the combination of two analytical tools: an elliptic regularity theory for a Neumann problem in divergence form with concentration-dependent coefficient (cf. (2.8)), and the regularity $L_{\operatorname{uloc}}^{4}([0,\infty);H^{2}(\Omega))$ for weak solutions, firstly discovered in [27]. Regarding the propagation of regularity, we will first revise the approach in [5]. By carefully examining the dependence of the constants appearing in the proof, our analysis shows that the correct differential inequality for $\Lambda$ reads as follows:
@@
for some positive constant $C$ independent of $\rho$, which differs from (1.15). Due to the factor $\frac{1}{\rho}$, the argument in [5] do not lead to a global bound for arbitrarily large initial datum. To overcome this issue, we will exploit a recent Grownall-type argument devised in [35, Lemma 2.1] (see Lemma 4.2 below). However, a direct application of this result leads to a global bound which grows double exponentially in time (4.25). Then, we propose a new approach to estimate the nonlinear term in (4.8). Once again, it relies on a careful use of the elliptic regularity theory for (2.8). Thanks to the regularity $\varphi\in L_{\operatorname{uloc}}^{4}([0,\infty);H^{2}(\Omega))$, the final differential inequality (4.27) fits into the classical framework of the uniform Gronwall lemma, thus implying the desired global estimate independent of time. Finally, the proof of the convergence of the weak solution towards a steady state of the Cahn-Hilliard equation is performed by following the well established method in the constant mobility case (e.g., [4]). This hinges upon the regularity of the weak solutions, the presence of a strict Lyapunov function, as well as the Łojasiewicz-Simon inequality.
@@ 2.1. Notation and function spaces
$\left\|f\right\|_{L^{2}(\Omega)}\leq C_{P}\left(\left\|\nabla f\right\|_{L^{2}(\Omega)}+|\overline{f}|\right)\!,\quad\forall\,f\in H^{1}(\Omega),$ (2.3)
@@
The solution operator $\mathcal{G}:H^{-1}_{(0)}(\Omega)\to H^{1}_{(0)}(\Omega)$ is defined as follows: for every $f\in H^{-1}_{(0)}(\Omega)$, $\mathcal{G}f\in H^{1}_{(0)}(\Omega)$ is the unique function satisfying
@@
where $F$ is the convex part of $\Psi$ as in (1.6). The functional $E_{0}$ is proper, convex and lower-semicontinuous. Owing to [39, Lemma 4.1], we deduce that $t\mapsto E_{0}(\varphi(t))$ is absolutely continuous on $[0,\infty)$ and
@@
almost everywhere on $(0,\infty)$. As a consequence, for any $T>0$, due to the boundedness of $F$ and $\varphi\in C([0,T];L^{2}(\Omega))$, the Lebesgue theorem entails that $\int_{\Omega}F(\varphi(\cdot))\ {\rm d}x\in C([0,T]),$ which in turn gives $\varphi\in C([0,T];H^{1}(\Omega))$. Now, taking $v=\mu$ in (1.20a) and exploiting the standard chain rule in $L^{2}(0,T;H^{1}(\Omega))\cap H^{1}(0,T;H^{1}(\Omega)^{\prime})$, we obtain
@@
for some constant $C$ depending on $E(\varphi_{0})$ and $m$. The latter implies $\nabla\mu\in L^{2}(0,\infty;L^{2}(\Omega))$. By (2.10) and (3.1), we have
@@
Next, multiplying (1.20b) by $-\Delta\varphi$, integrating over $\Omega$ and exploiting the boundary condition on $\varphi$, we find
@@
It is well known that the second term on the left-hand side is non-negative by the convexity of $F$ (see, e.g., [13, Appendix A]). Then, by (3.5), it follows that
@@
where the positive constant $C$ depends on $E(\varphi_{0})$ and $m$. Since $|\overline{\mu}|=\overline{\Psi^{\prime}(\varphi)},$ (3.12) and the above inequality imply that
@@
$\begin{cases}-\Delta\varphi+F^{\prime}(\varphi)=\mu^{*}\quad&\text{in }\Omega\\ \partial_{\textbf{{n}}}\varphi=0\quad&\text{on }\partial\Omega,\end{cases}$
@@
$\mu_{i}=-\mathcal{G}_{\varphi_{i}}\partial_{t}\varphi_{i}+\overline{\Psi^{\prime}(\varphi_{i})},\quad\text{ in }\Omega,\ \text{a.e. in }(0,\infty),\ i=1,2,$ (3.15)
where $\mu_{1}$ and $\mu_{2}$ are the chemical potentials corresponding to $\varphi_{1}$ and $\varphi_{2}$, respectively. We define $\Phi=\varphi_{1}-\varphi_{2}$, which solves
@@
We preliminary notice that $\overline{\Phi}\equiv 0$ by the conservation of mass. Multiplying the above equation by $\Phi$ and integrating over $\Omega$, we find
@@
Thanks to (3.15), we rewrite the above as
@@
almost everywhere in $(0,\infty)$. Recalling that $\varphi_{1}\in L^{\infty}(0,T;H^{1}(\Omega))\cap L^{4}(0,T;H^{2}(\Omega))$ with $\partial_{t}\varphi\in L^{2}(0,T;H^{-1}_{(0)}(\Omega))$, for any $T>0$, it follows by the elliptic regularity theory that
$\varphi_{1}^{\varepsilon}\in L^{\infty}(0,T;H^{3}(\Omega)),\quad\partial_{t}\varphi_{1}^{\varepsilon}\in L^{2}(0,T;H^{1}_{(0)}(\Omega)),$ (3.22)
and $\partial_{\textbf{{n}}}\varphi_{1}^{\varepsilon}=0$ on $\partial\Omega$ almost everywhere in $(0,T)$, for any $\varepsilon\in(0,1)$ and any $T>0$. In addition, recalling that $\varphi_{1}\in L^{\infty}(\Omega\times(0,T))$ with $|\varphi_{1}|<1$ almost everywhere in $\Omega\times(0,T)$, we also have
@@
$\int_{0}^{T}\left\|\varphi_{1}^{\varepsilon}(s)\right\|_{L^{2}(\Omega)}^{4}+\|\nabla\varphi_{1}^{\varepsilon}(s)\|_{L^{2}(\Omega)}^{4}\,{\rm d}s$ $\leq\int_{0}^{T}\|\varphi_{1}(s)\|_{L^{2}(\Omega)}^{4}+\|\nabla\varphi_{1}(s)\|_{L^{2}(\Omega)}^{4}\,{\rm d}s,$ (3.24)
@@
Owing to the regularity of $\varphi_{1}^{\varepsilon}$, we compute the gradient of (3.21) and test it by $-\nabla\Delta\varphi_{1}^{\varepsilon}$. Exploiting the Neumann boundary condition on $\varphi_{1}^{\varepsilon}$ and $\varphi_{1}$, we find
@@
$\varphi_{1}^{\varepsilon}\to\varphi_{1}\quad\text{strongly in }L^{4}(0,T;H^{2}(\Omega)),\quad\partial_{t}\varphi_{1}^{\varepsilon}\to\partial_{t}\varphi_{1}\quad\text{strongly in }L^{2}(0,T;H^{-1}_{(0)}(\Omega)).$ (3.28)
@@
$\left\|\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\right\|_{L^{\infty}(0,T;H^{1}_{(0)}(\Omega))}\leq C,$ (3.30)
@@
$\left\|\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\right\|_{H^{2}(\Omega)}\leq C\left(1+\|\varphi_{1}^{\varepsilon}\|_{H^{2}(\Omega)}\right)\!,$
@@
$\int_{0}^{T}\left\|\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi(s)\right\|_{H^{2}(\Omega)}^{4}\,{\rm d}s\leq C,$ (3.31)
@@
$\left\|\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\right\|_{W^{2,4}(\Omega)}\leq C\left(1+\left\|\varphi_{1}^{\varepsilon}\right\|_{H^{2}(\Omega)}^{\frac{3}{4}}\left\|\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\right\|_{H^{2}(\Omega)}^{\frac{3}{4}}\right)\!.$ (3.32)
@@
$\int_{0}^{T}\|\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi(s)\|_{W^{2,4}(\Omega)}^{\frac{8}{3}}\,{\rm d}s\leq CT+C\int_{0}^{T}\left\|\varphi_{1}^{\varepsilon}(s)\right\|_{H^{2}(\Omega)}^{4}+\left\|\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi(s)\right\|_{H^{2}(\Omega)}^{4}\,{\rm d}s\leq C,$ (3.33)
@@
$\int_{0}^{T}\left\|\nabla\left(\mathcal{G}_{\varphi_{1}}f-\mathcal{G}_{\varphi_{1}^{\varepsilon}}f\right)\right\|_{L^{2}(\Omega)}^{2}\,{\rm d}s$ $\leq\frac{1}{b_{m}}\int_{0}^{T}\left\|\left(b(\varphi_{1}^{\varepsilon})-b(\varphi_{1})\right)\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}f\right\|_{L^{2}(\Omega)}^{2}\,{\rm d}s$
$\leq\frac{1}{b_{m}}\int_{0}^{T}\left\|b(\varphi_{1}^{\varepsilon})-b(\varphi_{1})\right\|_{L^{\infty}(\Omega)}^{2}\left\|\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}f\right\|_{L^{2}(\Omega)}^{2}\,{\rm d}s.$
We aim to use the Lebesgue convergence theorem on the right-hand side. We observe from (2.10) that
$\left\|b(\varphi_{1}^{\varepsilon})-b(\varphi_{1})\right\|_{L^{\infty}(\Omega)}^{2}\left\|\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}f\right\|_{L^{2}(\Omega)}^{2}\leq\frac{4b_{M}}{b_{m}}\left\|\nabla\mathcal{G}f\right\|_{L^{2}(\Omega)}^{2},$
@@
$\left\|b(\varphi_{1}^{\varepsilon})-b(\varphi_{1})\right\|_{L^{\infty}(\Omega)}^{2}\leq C\left\|\varphi_{1}^{\varepsilon}-\varphi_{1}\right\|_{L^{\infty}(\Omega)}^{2}\leq C\left\|\varphi_{1}^{\varepsilon}-\varphi_{1}\right\|_{H^{2}(\Omega)}^{2}$
@@
$\left\|b(\varphi_{1}^{\varepsilon})-b(\varphi_{1})\right\|_{L^{\infty}(\Omega)}^{2}\left\|\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}f\right\|_{L^{2}(\Omega)}^{2}\leq C\|\varphi_{1}^{\varepsilon}-\varphi_{1}\|_{H^{2}(\Omega)}^{2}\|\nabla\mathcal{G}f\|_{L^{2}(\Omega)}^{2}\to 0$
@@
$\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}f\to\nabla\mathcal{G}_{\varphi_{1}}f\quad\text{strongly in }L^{2}(0,T;L^{2}(\Omega)).$